# Proposal for Improving Google A2A Protocol: Safeguarding Sensitive Data in Multi-Agent Systems

### v2

@@ 3.4 Lack of Transparency and User Consent
A2A lacks mechanisms that notify users about, or request consent for, the sharing of sensitive data with agents. “AI Agents Meet Blockchain” [32] emphasizes the importance of user transparency in decentralized agent-based environments. CVE-2024-44131 illustrates the consequences of bypassing such protections: malicious applications were able to circumvent Apple’s Transparency, Consent, and Control (TCC) framework to access sensitive data without user approval [5]. CWE-200 [9] codifies this as unauthorized exposure of sensitive information.
3.5 Potential Excessive Exposure of Data to Agents
@@
Agents in A2A ecosystems can access significantly more data than is necessary. “AI Agents Under Threat” [20] articulates how agent2agent data propagation can lead to unintended sharing of sensitive information. CVE-2023-41745 [3] and CVE-2022-45449 [2] document such exposures due to excessive privilege allocation or excessive data collection.
3.6 Risk of Data Disclosure to the Agent Itself
Moreover, the problem is not only that information is being disclosed to malicious actors, but that the very act of disclosing information to the agent itself is problematic, and that even well-meaning agents cannot be presumed to be safe from interference. Prompt injection attacks illustrate how malicious inputs can harvest classified information. “AI Agents Under Threat” [20] warns of this threat. CVE-2024-7042 [7] and CVE-2024-45989 [6] demonstrate how AI agents were exploited to perform unauthorized actions and exfiltrate sensitive data.
3.7 Concrete Example: Vacation Booking
Below is a usage scenario for the protocol without the proposed improvements. As reflected in Figure 2
  • The user requests an AI agent to book a vacation, including flights, hotel, and taxi, using his calendar, personal information and payment details.
  • The agent shares a payment token with the booking agent, valid for hours or days.
  • The agent accesses the user’s full calendar, including irrelevant personal details (e.g., medical appointments), and payment details.
Concern: The token could be used for additional payments, accidentally or intentionally, and calendar access exposes private information without explicit consent.
Figure 2: Vacation Booking Process Using An AI Agent
4 Proposed Enhancements to the Protocol
To address the identified issues, we propose seven enhancements to the A2A protocol, integrating research-backed solutions and aligning with advanced standards:
  1. 1.
    Short-Lived Tokens for Sensitive Operations
    Description: Mandate that tokens for sensitive operations, such as payments or ID transfers, have very short lifetimes (30 seconds to 5 minutes). Tokens expire after a single use or predefined time, requiring re-authentication for further actions.
    Rationale: As demonstrated in “AgNet” [25], short-lived tokens reduce the risk of unauthorized or repeated access. In our case, a payment token expiring quickly limits the window for leaks. “AI Agents Meet Blockchain” [32] supports this by noting that encryption and temporary records in blockchain systems enhance privacy, and our proposed tokens align with this principle.
  2. 2.
    Strong Customer Authentication (SCA) for Sensitive Transactions
    Description: Implement a mechanism requiring SCA, such as SMS codes, biometric verification, or bank login, for every sensitive transaction. This can be integrated into the task flow, ensuring user authentication before actions.
    Rationale: “AI Agents Meet Blockchain” [32] highlights that ZKPs enable secure authentication without exposing data, and in our case, SCA based on ZKPs or similar techniques ensures only verified users authorize sensitive actions. This aligns with PSD2 requirements for payment transactions, ensuring regulatory compliance.
  3. 3.
    More Granular Token Scopes
    Description: Extend OAuth 2.0 scopes to define precise permissions, such as “payment of \$1000 to Hotel X on Date Y” or “access to calendar availability only.” This prevents token use for unauthorized actions.
    Rationale: “Multi-Agent Security Tax” [42] notes that high granularity in permissions reduces unnecessary data exposure. In our case, precise scopes ensure agents access only required data, aligning with GDPR’s Minimum Necessary Disclosure principle.
  4. 4.
    Explicit User Consent Mechanism
    Description: Add a “consent” field to message or task structures, requiring agents to obtain explicit user approval before sharing sensitive data, specifying data type, purpose, and recipient.
    Rationale: “AI Agents Meet Blockchain” [32] emphasizes transparency as critical for trust in decentralized systems. In our case, explicit consent enhances user trust and ensures GDPR compliance, reducing the risk of unapproved data sharing.
  5. 5.
    Direct Data Transfer
    Description: Enable direct transfer of sensitive data between the user and service provider, bypassing intermediary agents. For example, in payments, the agent redirects the user to the bank for direct transfer, with user verification.
    Rationale: Inspired by “Multi-Agent Security Tax” [42]’s “active vaccines” to prevent malicious data spread, direct transfers avoid exposing data to intermediaries. “AI Agents Under Threat” [20] supports this by emphasizing the need to limit data propagation, and in our case, direct transfers prevent the exposure of sensitive data to the agent and reduce the risks of leakage.
  6. 6.
    Support for Multi-Transaction Approval
    Description: Allow a single user approval for a series of related transactions (e.g. flight, hotel, taxi payments), with tokens restricted to that series, short-lived, and SCA-verified.
    Rationale: This balances security and convenience, as recommended in “Multi-Agent Security Tax” [42] for efficiency-security trade-offs. In our case, it reduces repeated authentications while maintaining strict token restrictions. Additionally, this solution addresses the well-documented issue of ”consent fatigue,” where users, overwhelmed by frequent approval requests, may approve actions without proper scrutiny, increasing security risks. As highlighted in the Palo Alto Networks article [53] repeated requests can lead to user desensitization, potentially resulting in the approval of critical actions like write operations without adequate attention, a phenomenon similar to MFA fatigue. By consolidating approvals for related transactions, our solution minimizes the frequency of requests, thereby reducing the risk of exploitation by malicious actors while enhancing the user experience.
  7. 7.
    Compliance with Payment Standards
    Description: Ensure the protocol supports standards like PSD2, requiring SCA for certain transactions, by defining interfaces for payment agents that incorporate these requirements.
    Rationale: Compliance with regulations is critical for adoption, as noted in “AI Agents Meet Blockchain” [32]. Our proposal ensures A2A aligns with regulatory requirements, enhancing its reliability.
4.1 Application to the Example
Protocol Usage Scenario After The Proposed Improvements:
  1. 1.
    The user requests the AI agent to book a vacation, including flights, hotel, and taxi.
  2. 2.
    The agent requests permission to access calendar availability only, displaying: “I will share availability dates with the booking agent. Approve?” The user approves.
  3. 3.
    The agent finds a suitable booking and requests payment approval: “I will process \$1000 for Hotel X, \$500 for Flight Y, and \$50 for Taxi Z. Approve?” The user approves.
  4. 4.
    The agent requests a 5-minute token from the user’s bank, scoped to “payment for vacation X on Date Y.”
  5. 5.
    The bank issues a 5-minute token for the approved transaction series.
  6. 6.
    The agent shares the token with the booking agent.
  7. 7.
    The booking agent requests the bank to process payments using the token.
  8. 8.
    The bank prompts the user to verify the action via SMS code or biometric authentication.
  9. 9.
    The user approves, and payments are processed.
  10. 10.
    The token expires, preventing reuse.
Advantages:
  • Short-Lived Token: Prevents reuse or unauthorized access.
  • SCA: Ensures user approval for each payment.
  • Precise Scope: Token restricted to approved transactions.
  • Explicit Consent: The user is informed and approves the sharing of data.
  • Minimal Access: The agent accesses only availability, not the full calendar.
  • Direct Transfer: In order to protect sensitive information from agents, payment is sent straight from the bank to the booking agent.
4.2 Research Base Rationale for This Proposal
This proposal is based on previous research and enhances A2A’s security, privacy, and efficiency:
  • Enhanced Security: Short-lived tokens and SCA reduce leak risks, as recommended in [20, 32, 25].
  • Regulatory Compliance: SCA and PSD2 support ensure compliance, as emphasized in [32].
  • Transparency and Trust: Explicit consent increases user trust, per [32].
  • Flexibility and Efficiency: Multi-transaction approval balances security and convenience, as suggested in [42].
  • Common Standarts: The enhancements integrate with OAuth 2.0 and PSD2, as proposed in [25].
  • Exposure Prevention: Direct transfers minimize leak risks, according to [20, 42].
5 Evidence-Based Support for Proposed Enhancements
In order to confirm the technical rationale behind the proposed improvements to the Google A2A Protocol, this section provides a literature review of each of them. Each section focuses on one proposed improvement and summarises previous literature reviews showing that the improvements are not only technically motivated but also based on established research in areas such as authentication protocols, cryptography, distributed identity systems and access control. Together, these references provide empirical and architectural support for the adoption of the proposed measures.
5.1 Use of Short-Lived Access Tokens
One of the core enhancements we proposed to improve the security of Google’s A2A protocol is the mandatory use of short-lived access tokens. This design decision substantially reduces the window of opportunity for replay attacks and limits the utility of any compromised credential. Multiple studies affirm this principle. Teng (2023) introduced ActionID, a machine-to-machine authentication protocol that issues tokens tightly scoped in both action and time. The author notes that “the short-lived token limits the time window for action execution,” thereby reducing the effectiveness of credential theft [49]. Similarly, Ohwo et al. (2024) proposed a blockchain-based smart home access system in which “short-lived tokens [are] used to mitigate risks like replay attacks and user profiling” [40]. These tokens expire rapidly, neutralizing any intercepted credentials. Narajala et al. (2025) extended this logic to GenAI multi-agent systems, proposing a just-in-time registry that dynamically provisions tokens only for the duration of a tool invocation. This architecture, they argue, “minimizes the attack surface associated with persistent credentials by dynamically provisioning short-lived access tokens only when needed” [38]. Finally, Xiao et al. (2024) addressed the same vulnerability in the IoT domain, where resource-constrained devices are especially susceptible to token compromise. They developed MCU-Token, which binds a short-lived, hardware-derived token to each request, thereby making every token instance single-use and resistant to replay [51]. Together, these findings offer both conceptual and empirical support for implementing short-lived tokens in the A2A protocol as a fundamental safeguard against credential replay and session hijacking.
5.2 Strong Customer Authentication (SCA) for Sensitive Transactions
Implementing Strong Customer Authentication (SCA) for sensitive operations in the A2A protocol, such as payments and identity verification, significantly improves protection against impersonation and unauthorized transactions. Recent research underscores the feasibility and necessity of using modern, privacy-preserving SCA mechanisms such as zero-knowledge proofs (ZKPs), biometrics, and multi-factor authentication (MFA). For example, Neera et al. propose a mobile payment protocol that leverages ZKPs and identity-based signatures to verify user identity without revealing sensitive data while ensuring compliance with financial regulations such as PSD2 [39]. Their scheme guarantees cryptographically enforced SCA while maintaining user privacy, demonstrating that SCA can be both secure and regulatory compliant. Ahmad et al. present BAuth-ZKP, a blockchain-based MFA framework using smart contracts and ZKPs to authenticate users in smart city environments [15]. This design proves the viability of decentralized, privacy-respecting SCA mechanisms, directly applicable to agent-mediated A2A architectures. On the biometric front, Gernot and Rosenberger introduce a technique for generating one-time biometric templates, mitigating the risk of replay attacks by ensuring biometric credentials are valid only once [22]. This reinforces the ’inherence’ factor of SCA in a technically sound manner. Finally, Lyastani et al. empirically demonstrate that while 2FA dramatically improves account security, inconsistent or poorly designed SCA flows reduce adoption [23]. Thus, their findings support designing usable, transparent SCA processes for sensitive actions in A2A. Collectively, these studies provide strong empirical and architectural justification for integrating SCA into the A2A protocol, ensuring that sensitive transactions are approved only by verified users under secure and user-friendly conditions.
5.3 More Granular Token Scopes
Enforcing fine-grained token scopes in the A2A protocol is essential for upholding the principle of least privilege and minimizing the exposure of sensitive resources during agent-to-agent transactions. The risks of overly broad permissions in bearer tokens are well-documented in recent research. Cao et al. propose a stateful, least-privilege authorization model that allows client-side applications to dynamically constrain the scope of OAuth tokens using WebAssembly-based privilege attenuation policies [18]. Their system empowers developers to explicitly encode minimal access rights per session, enabling secure delegation without overprovisioning. This aligns directly with the A2A context, where agents must act within strict permission boundaries to avoid inadvertent data access. Complementarily, South et al. extend OAuth 2.0 and OpenID Connect for authenticated agent delegation, introducing agent-specific credentials with precise, auditable scope limitations [47]. Their framework demonstrates how user intent can be translated into tightly scoped permissions, ensuring that AI agents operate only within authorized domains. These works collectively support the need for scope-constrained access tokens as a foundational safeguard in agent-mediated architectures. Dimova et al. further substantiate this by showing that 18.5% of OAuth deployments on the web request unnecessary scopes, violating the GDPR’s minimum necessary data principle [21]. Kaltenböck et al. reinforce this position by embedding scope-aware policies within a Zero Trust single sign-on framework, emphasizing explicit scope definitions to limit token capabilities at authentication time [31]. Altogether, the literature affirms that fine-grained token scopes are not only technically viable but also indispensable for ensuring data minimization, secure delegation, and regulatory compliance in distributed, AI-driven authentication systems.
5.4 Explicit User Consent Mechanism
Embedding an explicit user consent mechanism within the A2A protocol is critical for aligning with privacy regulations such as GDPR and for building user trust in agent-mediated systems. Recent research emphasizes that consent must be freely given, informed, specific, and revocable throughout the data lifecycle. Merlec et al. propose a blockchain-based dynamic consent management system, where users can grant, audit, or withdraw their consent via smart contracts stored on a tamper-proof ledger [36]. This architecture guarantees accountability, traceability, and user autonomy in data-sharing environments. Complementarily, Khalid et al. formalize the security and privacy requirements of such systems, proposing the integration of zero-knowledge proofs and cryptographic primitives to ensure that consent is provable, minimal, and compliant by design [33]. Their work outlines how systems can enforce consent boundaries while maintaining confidentiality. In a multi-agent context, Xu et al. demonstrate that autonomous privacy agents can enforce user-defined consent policies, making decisions that reflect GDPR principles such as data minimization and informed processing [52]. These agents act only within verified constraints, providing technical assurance that consent decisions are respected. Finally, Pathmabandu et al. introduce a consent management engine that enables granular, real-time visibility into data collection within IoT systems [41]. Their engine offers digital nudging and fine-grained control, reinforcing the user’s right to control their personal data. Collectively, these studies support the integration of explicit, technically enforceable consent mechanisms into A2A-style protocols, ensuring that sensitive data is shared only with informed user approval, and that such actions remain transparent and auditable.
5.5 Direct Data Transfer
In the evolving landscape of multi-agent systems, ensuring the security of sensitive data during interactions between users and service providers is paramount. Traditional multi-agent architectures often involve intermediaries that can pose significant security risks, such as data leakage or unauthorized access. To address these concerns, the proposed enhancement advocates for direct data transfer, where sensitive information is exchanged directly between the user and the service provider, bypassing intermediary agents. This approach is supported by recent research in multi-agent security. Firstly, [42] introduces the concept of ”active vaccines” to prevent the spread of malicious prompts in multi-agent systems, underscoring the need to minimize intermediary involvement to reduce systemic vulnerabilities. Similarly, [20] highlights the unpredictability of multi-step user inputs and the complexity of internal executions in AI agents, emphasizing the necessity of limiting data propagation to trusted entities. Furthermore, [19] discusses the open challenges in securing systems of interacting AI agents, particularly the threats arising from free-form interactions and network effects that can amplify security breaches, which direct data transfer mitigates by reducing intermediaries. Lastly, [47] presents a framework for authenticated delegation and authorized AI agents, which can be extended to support direct data transfer by ensuring secure, intermediary-free communication channels. On the basis of these studies, it can be concluded that direct transmission of data increases the security of multi-agent systems by reducing reliance on intermediaries and thus minimising the risks of data exposure and unauthorised access. This improvement is essential for applications involving sensitive data such as payments and personal data, where security and privacy are paramount.
5.6 Support for Multi-Transaction Approval
In the area of multi-agent systems and financial technology, securing financial transactions while preserving user comfort is a key challenge, especially for sensitive data such as payment details. The proposed enhancement, ’Multi-Transaction Approval,állows for a single authorization for a series of transactions using limited-time tokens validated by strong customer authentication. This approach balances security and usability by reducing the need for multiple authentication while maintaining strict security measures, thus increasing the effectiveness of multi-step workflow in agent-based systems. Recent research on the security of financial technology supports this improvement. For instance, [34] underscore the pivotal role of multi-factor authentication (MFA) in securing mobile financial transactions, advocating its use to mitigate fraud risks, which aligns with the SCA requirements of the proposed enhancement. Similarly, [50] provide a systematic review of MFA in digital payment systems, noting that grouping transactions under a single secure authentication session is feasible with robust mechanisms, supporting the enhancement’s design. Furthermore, [13] propose a framework integrating MFA with machine learning to secure online financial transactions, adaptable to multi-transaction approval by ensuring SCA verification for each transaction in a series and employing anomaly detection to enhance security. Additionally, [16] develop an MFA algorithm for mobile money applications, combining multiple authentication factors to secure transactions, which can be extended to support a single approval for multiple transactions as proposed. In conclusion, the “Multi-Transaction Approval” enhancement is robustly supported by current research in financial technology security, emphasizing advanced MFA techniques and fraud detection methods. By implementing this enhancement, the A2A protocol can offer a secure and user-friendly approach to managing multiple related transactions, reducing user friction while adhering to high security standards.
5.7 Compliance with Payment Standards
In the area of multi-agent systems and financial technology, compliance with regulatory standards such as PSD2 is essential to ensure the safe and secure processing of transactions. The proposed enhancement, “Compliance with Payment Standards”, integrates these requirements into Google’s A2A protocol by defining interfaces for payment agents that include strong customer authentication and other regulatory requirements. This approach increases the credibility and readiness of the Protocol for implementation in the real world by reputable entities. Recent research supports the feasibility of such improvements. For instance, [32] highlight how blockchain technology can ensure compliance through tamper-proof documentation, aligning with PSD2’s demands for transparency and security in financial transactions [32]. Similarly, [17] demonstrate that artificial intelligence can bolster regulatory compliance in the financial sector by leveraging machine learning to monitor and prevent breaches, a principle applicable to multi-agent systems to enforce PSD2 standards. Lastly, [46] address the broader compliance landscape for AI systems, focusing on the EU’s AI Act and data set compliance, which reinforces the importance of embedding regulatory adherence into AI-driven systems, including those involving multi-agent interactions [46]. By integrating these insights, the enhancement positions the A2A protocol as a robust framework for secure, compliant financial transactions in multi-agent environments.
6 Prompt Injection in LLM-Based Agents
Prompt injection attacks are a significant concern in the domain of artificial intelligence security, particularly with the widespread deployment of large language models (LLMs) in applications ranging from automated customer support to complex data processing systems. These attacks involve the deliberate insertion of malicious or deceptive text inputs designed to manipulate an AI model’s behavior, potentially leading to the disclosure of sensitive information, the generation of harmful outputs, or the undermining of the system’s intended purpose. The vulnerability arises from the inherent difficulty AI systems face in differentiating between legitimate user inputs and crafted malicious prompts, a challenge that becomes increasingly significant as AI integrates into sensitive operational contexts.
Already in 2022, as LLMs began their global proliferation, researchers identified the risks posed by malicious prompt manipulation. Perez et al. [43] demonstrated that carefully designed input could override predefined model instructions, forcing the model to ignore prior constraints, disclose sensitive data, or behave in unintentional ways. Although initial defenses have since improved, new and more sophisticated methods continue to evolve. In 2023, Liu et al. [35] revealed that 31 of 36 LLM-integrated applications tested, including productivity and chat tools, remained vulnerable to prompt injection, even in black-box scenarios. Their work underscores that the threat landscape is not only persistent, but also increasingly complex.
Figure 3: Prompt Injection Proccess Via N8N Platform - Agent 1
To investigate these threats in practice, we conducted an experiment using the N8N platform, a workflow automation tool that facilitates the creation of AI agents. In this experiment, a first AI agent, as shown in Figure 3, was constructed with explicit instructions not to reveal private information critical to its function, simulating a scenario where confidentiality is paramount. Subsequently, a second AI agent, as shown in Figure 4, was developed with the sole purpose of extracting this withheld information from the first agent. Employing various manipulative techniques and prompt injection strategies, the second agent successfully elicited the sensitive data, without relying on advanced tactics, thereby exposing the fragility of AI systems to such attacks and highlighting the practical risks of entrusting them with confidential information.

### v3

@@ IV-D Lack of Transparency and User Consent
Problem: A2A lacks mechanisms for informing users or obtaining consent before sensitive data is shared between agents. As illustrated in Figure 3, the agent’s access to unrelated calendar entries (e.g., medical appointments) exemplifies the lack of explicit user consent in sensitive data sharing. Research emphasizes that user transparency is essential in decentralized environments [8]. For example, CVE-2024-44131 [46] describes how malicious applications bypassed Apple’s Transparency Consent and Control (TCC) framework to access private data without user approval. CWE-200 [47] codifies this as unauthorized exposure of sensitive information.
Proposed Solution: To address this, we propose an explicit user consent mechanism requiring agents to include explicit, user-approved consent metadata before transmitting sensitive information. This includes specifying data type, purpose, and recipient. Transparent consent flows are essential in decentralized systems. They improve user trust and ensure compliance with GDPR and similar regulations. Explicit, verifiable user consent strengthens privacy and legal compliance. Recent research emphasizes that consent must be freely given, informed, specific, and revocable throughout the data lifecycle. Merlec et al. propose a blockchain-based dynamic consent management system, where users can grant, audit, or withdraw their consent via smart contracts stored on a tamper-proof ledger [48]. This architecture guarantees accountability, traceability, and user autonomy in data-sharing environments. Complementarily, Khalid et al. formalize the security and privacy requirements of such systems, proposing the integration of zero-knowledge proofs and cryptographic primitives to ensure that consent is provable, minimal, and compliant by design [49]. Their work outlines how systems can enforce consent boundaries while maintaining confidentiality. In a multi-agent context, Xu et al. demonstrate that autonomous privacy agents can enforce user-defined consent policies, making decisions that reflect GDPR principles such as data minimization and informed processing [50]. These agents act only within verified constraints, providing technical assurance that consent decisions are respected. Finally, Pathmabandu et al. introduce a consent management engine that enables granular, real-time visibility into data collection within IoT systems [51]. Their engine offers digital nudging and fine-grained control, reinforcing the user’s right to control their personal data. Collectively, these studies support the integration of explicit, technically enforceable consent mechanisms into A2A-style protocols, ensuring that sensitive data is shared only with informed user approval, and that such actions remain transparent and auditable.
IV-E Potential Excessive Exposure of Data to Agents
@@
Problem: Figure 3 also demonstrates how excessive agent access to personal data, such as entire calendar contents, can lead to unnecessary exposure. Agents in A2A ecosystems can access significantly more data than is necessary. “AI Agents Under Threat” [7] articulates how agent-to-agent data propagation can lead to unintended sharing of sensitive information. CVE-2023-41745 [52] and CVE-2022-45449 [53] document such exposures due to excessive privilege allocation or excessive data collection.
Proposed Solution: A suitable countermeasure is direct data transfer flows in which sensitive information (e.g., payments or ID documents) is sent directly between the user and the service provider, bypassing intermediary agents. This approach is supported by recent research in multi-agent security. Firstly, [13] introduces the concept of ”active vaccines” to prevent the spread of malicious prompts in multi-agent systems, underscoring the need to minimize intermediary involvement to reduce systemic vulnerabilities. Similarly, [7] highlights the unpredictability of multi-step user inputs and the complexity of internal executions in AI agents, emphasizing the necessity of limiting data propagation to trusted entities. Furthermore, [54] discusses the open challenges in securing systems of interacting AI agents, particularly the threats arising from free-form interactions and network effects that can amplify security breaches, which direct data transfer mitigates by reducing intermediaries. Lastly, [43] presents a framework for authenticated delegation and authorized AI agents, which can be extended to support direct data transfer by ensuring secure, intermediary-free communication channels. On the basis of these studies, it can be concluded that direct transmission of data increases the security of multi-agent systems by reducing reliance on intermediaries and thus minimising the risks of data exposure and unauthorised access. This improvement is essential for applications involving sensitive data such as payments and personal data, where security and privacy are paramount.
IV-F Risk of Data Disclosure to the Agent Itself
Problem: Even when agents are not malicious, granting them access to sensitive data remains risky. As depicted in Figure 3, the agent gains direct access to sensitive personal details, which heightens the risk of unintended disclosure even without malicious intent. Prompt injection attacks demonstrate how adversarial inputs can manipulate AI models to reveal confidential content [7]. CVE-2024-7042 [55] and CVE-2024-45989 [56] demonstrate how AI agents were exploited to perform unauthorized actions and exfiltrate sensitive data.
Proposed Solution: To mitigate this risk, we advocate minimizing agents access to sensitive data whenever feasible. This approach reduces the potential attack surface while preserving core functionality. A deeper exploration of this principle, including experimental validation and architectural considerations, is provided in subsequent sections.
IV-G Consent Fatigue in Multi-Transaction Workflows
Problem: In agent-mediated environments, users are often asked to repeatedly authenticate or approve actions within related operations. This can cause consent fatigue, where users grow desensitized to security prompts, increasing the risk of errors or inadvertent approvals. Such fatigue undermines usability and security, adding friction without real risk reduction.
Proposed Solution: To address this, we propose allowing users to approve a bundled set of related transactions (e.g., flight, hotel, taxi) in a single consent flow, using scoped, short-lived, SCA-verified tokens. This approach balances user experience and security, as grouped approval reduces repeated prompts and mitigates consent fatigue, a well-known issue in security usability studies [57, 13]. Recent research on the security of financial technology supports this improvement. For instance, [58] underscore the pivotal role of multi-factor authentication (MFA) in securing mobile financial transactions, advocating its use to mitigate fraud risks, which aligns with the SCA requirements of the proposed enhancement. Similarly, [59] provide a systematic review of MFA in digital payment systems, noting that grouping transactions under a single secure authentication session is feasible with robust mechanisms, supporting the enhancement’s design. Furthermore, [60] propose a framework integrating MFA with machine learning to secure online financial transactions, adaptable to multi-transaction approval by ensuring SCA verification for each transaction in a series and employing anomaly detection to enhance security. Additionally, [61] develop an MFA algorithm for mobile money applications, combining multiple authentication factors to secure transactions, which can be extended to support a single approval for multiple transactions as proposed. In conclusion, the “Multi-Transaction Approval” enhancement is robustly supported by current research in financial technology security, emphasizing advanced MFA techniques and fraud detection methods. By implementing this enhancement, the A2A protocol can offer a secure and user-friendly approach to managing multiple related transactions, reducing user friction while adhering to high security standards.
IV-H Regulatory Compliance Gaps
Problem: As agent ecosystems expand into sensitive domains such as finance and identity management, regulatory adherence is essential. The current A2A protocol lacks mechanisms to meet standards like PSD2, risking user trust, compliance, and enterprise adoption. Without support for secure authentication, transaction logging, and consent auditing, the protocol may fall short of legal and industry expectations.
Proposed Solution: We propose integrating A2A with regulations such as PSD2 by incorporating authentication, logging, and mandatory controls in agent interfaces. This approach increases the credibility and readiness of the Protocol for implementation in the real world by reputable entities. Recent research supports the feasibility of such improvements. For instance, [8] highlight how blockchain technology can ensure compliance through tamper-proof documentation, aligning with PSD2’s demands for transparency and security in financial transactions. Similarly, [62] demonstrate that artificial intelligence can bolster regulatory compliance in the financial sector by leveraging machine learning to monitor and prevent breaches, a principle applicable to multi-agent systems to enforce PSD2 standards. Lastly, [63] address the broader compliance landscape for AI systems, focusing on the EU’s AI Act and data set compliance, which reinforces the importance of embedding regulatory adherence into AI-driven systems, including those involving multi-agent interactions. By integrating these insights, the enhancement positions the A2A protocol as a robust framework for secure, compliant financial transactions in multi-agent environments.
IV-I Illustrative Secure Vacation Booking Flow
To demonstrate the practical application of the proposed enhancements, consider a secure vacation booking scenario that addresses the identified vulnerabilities. The user initiates a request for an agent to coordinate flights, hotel, and taxi reservations. Rather than granting broad calendar access, the agent requests access solely to availability data, which the user explicitly approves. Upon identifying suitable offers, the agent prompts the user for a bundled, SCA-verified consent covering all related transactions. The user’s bank then issues a short-lived token scoped to the approved amounts. This token is securely transmitted to the booking agent and is valid for a single use only. The bank completes the payment process following strong customer authentication, and the token expires immediately upon successful transaction completion. This flow illustrates how integrated safeguards, including granular permissions, short-lived credentials, explicit consent, and regulatory-aligned controls, can collectively mitigate risks while preserving usability and operational efficiency.
V Explicit User Consent Orchestration via USER_CONSENT_REQUIRED
Among the enhancements proposed in this paper, we chose to demonstrate the explicit consent mechanism due to its conceptual clarity, low integration barrier, and direct impact on regulatory compliance. The following code and design pattern illustrate how this enhancement can be implemented directly into the A2A protocol.
A new enumeration member is introduced in the TaskState enum:  [^1] A full code sample is available at: https://github.com/yedidel/A2A
USER_CONSENT_REQUIRED = "user-consent-required"
This addition extends the protocol’s control over task execution. Unlike INPUT_REQUIRED, which signals a need for additional data, USER_CONSENT_REQUIRED pauses execution until the end-user provides affirmative consent. For example, before processing a payment or releasing personal information, agents must now request user approval explicitly. As depicted in the diagram shown in Figure 4.
This new state enables precise front-end behavior. When encountered, the UI prompts the user to approve the requested action. Upon approval, the task transitions to IN_PROGRESS, continuing as normal. This pattern ensures user awareness and strengthens trust.
Moreover, the mechanism supports auditability and regulatory alignment. By recording the explicit consent event, the system can demonstrate compliance with GDPR, PSD2, and similar frameworks. This enhancement is therefore a concrete instantiation of both technical and ethical design principles outlined earlier in this paper.
Figure 4: Booking process flow that pauses at sensitive actions until explicit user approval is obtained using USER_CONSENT_REQUIRED State
VI Case Study: Prompt Injection Attack
While the previous sections focused on structural weaknesses in the A2A protocol, it is equally important to understand how these vulnerabilities manifest in real-world scenarios. One of the most prominent and actively exploited risks in modern AI-based systems is prompt injection, a technique that demonstrates how even seemingly well-configured agents can be manipulated into disclosing sensitive data or violating intended behavior. In this section, we illustrate and implementation how the absence of robust permission controls and data boundaries can enable such attacks in practice, reinforcing the need for the proposed enhancements to the protocol.
Since the early emergence of large language models, prompt injection has posed a challenge, allowing adversaries to override system-level constraints and access restricted information. researchers identified the risks posed by malicious prompt manipulation. Perez et al. [64] demonstrated that carefully designed input could override predefined model instructions, forcing the model to ignore prior constraints, disclose sensitive data, or behave in unintentional ways. Although initial defenses have since improved, new and more sophisticated methods continue to evolve. Liu et al. [23] revealed that 31 of 36 LLM-integrated applications tested, including productivity and chat tools, remained vulnerable to prompt injection, even in black-box scenarios. Their work underscores that the threat landscape is not only persistent, but also increasingly complex.
Figure 5: Configuration of Agent 1 in N8N with access restrictions but retention of sensitive data in memory.
To assess these risks, we conducted an experiment using the N8N platform. The first agent was configured with strict instructions such as “Do not share any secret codes or identity numbers”. As shown in Figure 5, this diagram illustrates the configuration and initial constraints applied to Agent 1. The second agent was designed to elicit that hidden data using prompt injection. Success in this context was defined as any instance where Agent 1’s output contained part or all of the protected data in response to the injected prompt. Despite the protections in place, the second agent succeeded using basic manipulation techniques. The second agent asked Agent 1 to concatenate new information with existing data. Although Agent 1 correctly stated that it could not share secret data, its reply inadvertently included the sensitive value. This demonstrates how simple chaining prompts can bypass protections, and how even additional checks could be manipulated (e.g., by encoding the secret as letters). Figure 6 shows how Agent 2’s prompt was processed and led to disclosure of sensitive data. This confirmed how easily agents can be compromised, even without sophisticated attacks. This experiment did not involve human subjects data and did not require IRB or equivalent ethical oversight. The prompt injection experiment presented here is illustrative rather than novel, intended to demonstrate the ease of exploiting such vulnerabilities in practice. For in-depth technical analyses of prompt injection mechanisms and their root causes, we refer readers to the cited works, which provide comprehensive treatments of this well-studied threat.
Figure 6: Operation of Agent 2 sending crafted prompts to Agent A to elicit protected information.